OpenStudy (anonymous):

what is the radius of the circle: (x-3)^2+(y+2)^2=14

OpenStudy (anonymous):

\[r= \sqrt {14}\]

OpenStudy (anonymous):

just compare it from standard equation of circle that is (x-a)^2+(x-b)^2=R^2

OpenStudy (anonymous):

here you go x^2 + y^2 = r^2. Since your equation says (x – 3)^2 + (y + 2)^2 = 14, 14 is the radius squared. Therefore, the radius of the circle for the second equation is the square root of 14, or 3.74165739.

OpenStudy (anonymous):

Thanks girl! I've got so many problems I'm trying to figure out! lol

OpenStudy (anonymous):

lol np